### v1

@@ 4.1 Algorithm framework
For $\psi\in\bigoplus_{i=1}^{q}M_{h}(\lambda_{i})$ with $\|\psi\|_{a}=1$ and the orthonormal basis $\{v^{h}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})$ with $a(v^{h}_{ij},v^{h}_{kl})=\delta_{ik}\delta_{jl}$, there exists $\{\alpha_{ij}\}$ satisfying $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\alpha_{ij}^{2}=1$ such that $\psi=\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\alpha_{ij}v^{h}_{ij}.$ It holds that
(17) $\operatorname{dist}_{a}\left(\psi,U_{n_{0}+1}\right)=\left\|\left(\operatorname{I}-\mathcal{P}_{U_{n_{0}+1}}\right)\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\alpha_{ij}v^{h}_{ij}\right\|_{a}$
$\leqslant$ $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}|\alpha_{ij}|\left\|\left(\operatorname{I}-\mathcal{P}_{U_{n_{0}+1}}\right)v^{h}_{ij}\right\|_{a}\leqslant\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\left\|\left(\operatorname{I}-\mathcal{P}_{U_{n_{0}+1}}\right)v^{h}_{ij}\right\|_{a}^{2}}$
$=$ $\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\operatorname{dist}_{a}^{2}\left(v^{h}_{ij},U_{n_{0}+1}\right)}\leqslant\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\operatorname{dist}_{a}^{2}\left(M_{h}(\lambda_{i}),U_{n_{0}+1}\right)}$
$\leqslant$ $\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\operatorname{dist}_{a}^{2}\left(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)}\right)}\leqslant\sqrt{N}\max_{1\leqslant i\leqslant q}\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)}),$
@@
(18) $\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n_{0}+1}\right)\leqslant\sqrt{N}\max_{1\leqslant i\leqslant q}\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)}).$
Moreover, there holds $\lambda^{h}_{ij}<\lambda_{i+1}$ due to $\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n_{0}+1}\right)\ll 1$ for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$. Hence, the error estimate for Algorithm 1 with solving a small eigenvalue problem in the $n_{0}$-th iteration follows from Eq. 18, Proposition 2.3 and Theorem 3.7 applied to Eq. 14 when $n=n_{0}$.
We will see in the next subsection that the requirements $\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)})\ll 1(i=1,\ldots,q)$ can be achieved. We understand that one of the main cost in Algorithm 1 is made by step $4$: the generation and solution of the $N$-dimensional eigenvalue problem. Fortunately, we see from the numerical experiments in [9, 22] that the resulting matrix is almost diagonal: the non-diagonal entries are very small and the computational cost is not very expensive.
4.2 Shifted-inverse based ParO algorithm
In this subsection, we shall study the ParO algorithm for solving the clustered eigenvalue problem when the shifted-inverse approach is applied to update each orbital. The shifted-inverse based ParO algorithm (Algorithm 3), which solves a small scale eigenvalue problem in each iteration to update the shift parameters, has been proposed in [9, 22]. To show the convergence, we first consider a simplified version (Algorithm 2) which fixes the shift parameters and does not carry out the steps of solving projected eigenvalue problems in iterations. Based on the numerical analysis of the simplified version, we then prove that the approximations produced by Algorithm 3 converge rapidly.
We set the shift parameter as any convex combination of $\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}$ denoted by
In our discussion, we assume that the shift parameters are always not equal to the eigenvalues of Eq. 3 in the calculation process. Otherwise, we continue the iterative process on other orbitals while keeping the orbitals unchanged.
Then the shifted-inverse approach $\mathcal{F}_{n}$ writes: for $U_{n}^{(i)}=\operatorname{span}\left\{u_{i1}^{(n)},\ldots,u_{id_{i}}^{(n)}\right\},$ $\mathcal{F}_{n}^{(i)}:=\mathcal{F}_{n}|_{U_{n}^{(i)}}:U_{n}^{(i)}\rightarrow U_{n+1/2}^{(i)}$ with $u_{ij}^{(n+1/2)}=\mathcal{F}_{n}^{(i)}u_{ij}^{(n)}$ for $i=1,\ldots,q,j=1,2,\ldots,d_{i}$ satisfying
The simplified shifted-inverse based ParO algorithm is stated as Algorithm 2. Compared with Algorithm 1, step 4 is no longer carried out in this simplified version.
Algorithm 2 Simplified shifted-inverse based ParO algorithm
  1. 1.
    Given a finite-dimensional space $V^{h}$ and $tol>0$, provide and cluster initial data by Eq. 12, i.e., $\left\{\left(\lambda_{ij}^{(0)},u_{ij}^{(0)}\right)\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}\subset\mathbb{R}\times V^{h}$ with $b\left(u_{ij}^{(0)},u_{kl}^{(0)}\right)=\delta_{ik}\delta_{jl}$. Set $\bar{\lambda}_{i}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}\right)$ and let $n=0$.
  2. 2.
  3. 3.
    Set $u_{ij}^{(n+1)}=\frac{u_{ij}^{(n+1/2)}}{\left\|u_{ij}^{(n+1/2)}\right\|_{b}}$. If $\frac{\|u_{ij}^{(n+1)}-u_{ij}^{(n)}\|_{b}}{\|u_{ij}^{(n)}\|_{b}}>tol$, let $n=n+1$ and go to $2$.
If the initial guesses approximate the exact eigenvalues good enough, then the source problems Eq. 19 will be ill-conditioned. We mention that there are approaches to deal with ill-conditioned systems (see e.g., [3, 5, 6, 25, 27]). Indeed, it is quite difficult to solve these ill-conditioned systems well, which will be discussed in our other work. Here we assume that such systems can be well solved.
Proposition 4.3.
If $U^{(i)}_{n+1/2}$ is obtained by Algorithm 2, then
Proof 4.4.
@@ 4.2 Shifted-inverse based ParO algorithm
With the help of Proposition 4.3, we obtain convergence of eigenspace approximations produced by Algorithm 2.
Theorem 4.5.
Assume that
(21) $\operatorname{dim}\left(U_{0}^{(i)}\right)=d_{i},\quad\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{0}^{(i)}\right)<1\quad\forall i=1,2,\ldots,q.$
If $U_{n+1/2}^{(i)}$ is produced by Algorithm 2, then
$\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}\right)\leqslant\varepsilon_{n+1}=\frac{\delta_{0}\varepsilon_{n}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{n}^{2})+\delta_{0}^{2}\varepsilon^{2}_{n}}},\quad\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}}=\frac{\delta_{0}}{g-\delta_{0}}.$
Proof 4.6.
@@
A simple calculation and
$a(\mathcal{P}_{M_{h}(\lambda_{i})}v_{i}^{(0)},v)=a(v_{i}^{(0)},v),\quad v\in M_{h}(\lambda_{i}),$
@@
We obtain from Lemma 2.2 and Eq. 21 that there exists $\varepsilon_{0}\in(0,1)$ such that
$\varepsilon_{0}\geqslant$ $\operatorname{dist}_{a}\left(U_{0}^{(i)},M_{h}(\lambda_{i})\right)\geqslant\operatorname{dist}_{a}\left(v_{i}^{(0)},M_{h}(\lambda_{i})\right)$
$=$ $\operatorname{dist}_{a}\left(v_{i}^{(0)},\mathcal{P}_{M_{h}(\lambda_{i})}v_{i}^{(0)}\right)=\frac{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\alpha_{rt}^{(i)}u^{h}_{rt}-\sum_{t=1}^{d_{i}}\alpha_{it}^{(i)}u^{h}_{it}\right\|_{a}}{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\alpha_{rt}^{(i)}u^{h}_{rt}\right\|_{a}},$
which yields,
@@
(24) $\operatorname{dist}_{a}\left(v_{i}^{(1/2)},M_{h}(\lambda_{i})\right)=$ $\frac{\left\|v_{i}^{(1/2)}-\mathcal{P}_{M_{h}(\lambda_{i})}v_{i}^{(1/2)}\right\|_{a}}{\left\|v_{i}^{(1/2)}\right\|_{a}}$
@@
$\leqslant$ $\frac{\delta_{0}\varepsilon_{0}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{0}^{2})+\delta_{0}^{2}\varepsilon^{2}_{0}}}\triangleq\varepsilon_{1},\quad i=1,\ldots,q.$
@@
$\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{1/2}^{(i)}\right)\leqslant\varepsilon_{1},\quad i=1,2,\ldots,q.$
@@
$\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}\right)\leqslant\varepsilon_{n+1}=\frac{\delta_{0}\varepsilon_{n}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{n}^{2})+\delta_{0}^{2}\varepsilon^{2}_{n}}},\quad\forall n\in\mathbb{N}.$
@@
Theorem 4.5 shows that convergence of Algorithm 2 does not require sufficiently accurate initial guesses. Eq. 20 ensures that the shift parameter is closer to the eigenvalue being approximated. Indeed, Eq. 20 can be satisfied under the assumption that the finite dimensional discretization Eq. 3 approximates Eq. 1 not “too badly”. Eq. 21 guarantees that the dimension of the approximated subspace is preserved. When $N=1$, Eq. 25 may be reviewed as the classical shift-inverse convergence result.
Algorithm 3 Shifted-inverse based ParO algorithm
  1. 1.
    Given a finite-dimensional space $V^{h}$ and $tol>0$, provide and cluster initial data by Eq. 12, i.e., $\left\{\left(\lambda_{ij}^{(0)},u_{ij}^{(0)}\right)\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}\subset\mathbb{R}\times V^{h}$ with $b\left(u_{ij}^{(0)},u_{kl}^{(0)}\right)=\delta_{ik}\delta_{jl}$. Set $\bar{\lambda}_{i}^{(0)}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}\right)$ and let $n=0$.
  2. 2.
  3. 3.
    Construct $U_{n+1}=\operatorname{span}\left\{u_{11}^{(n+1/2)},\ldots,u_{1d_{1}}^{(n+1/2)},\ldots,u_{qd_{q}}^{(n+1/2)}\right\}$.
  4. 4.
    Solve an eigenvalue problem: find $(\lambda^{(n+1)},u^{(n+1)})\in\mathbb{R}\times U_{n+1}$ satisfying
  5. 5.
    If $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\left|\lambda_{ij}^{(n+1)}-\lambda_{ij}^{(n)}\right|>tol$, set $\bar{\lambda}_{i}^{(n+1)}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(n+1)}\right\}_{j=1}^{d_{i}}\right)$, $n=n+1$ and go to $2$.
As mentioned above, we assume that the shift parameters are always not equal to the eigenvalues of Eq. 3 in the calculation process. If there are cases where some orbitals are very well approximated, while other orbitals have not yet converged, then we continue the iterative process on the non-converged orbitals while keeping the well approximated orbitals unchanged.
The following theorem tells the convergence of Algorithm 3.
Theorem 4.7.
(28) $\operatorname{dist}_{a}\left(u^{h,o,0}_{ij},u_{ij}^{(0)}\right)$ $\leqslant\varepsilon_{0},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
$\operatorname{dist}_{a}\left(u^{h,o,n+1}_{ij},u_{ij}^{(n+1)}\right)$ $\leqslant\varepsilon_{n+1}=\frac{C\sqrt{DN}\left(\gamma+\zeta_{n}\right)\varepsilon_{n}}{\sqrt{(g-\gamma-\zeta_{n})^{2}(1-D\varepsilon_{n}^{2})+D\left(\gamma+\zeta_{n}\right)^{2}\varepsilon^{2}_{n}}},$
$|\lambda^{h}_{ij}-\lambda_{ij}^{(n+1)}|$ $\leqslant\zeta_{n+1}:=\frac{\lambda^{h}_{q+1,1}}{C^{2}}\varepsilon_{n+1}^{2},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $D:=\max_{1\leqslant i\leqslant q}d_{i}$ and $C$ is a constant that comes from the application of Theorem 3.7, and is independent of $U_{n}(n=0,1,2,\ldots)$, and
(30) $\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}}=\frac{C\sqrt{DN}\gamma}{g-\gamma}.$
Proof 4.8.
@@
Since $\{u^{h,o,0}_{ij}\}_{i=1}^{d_{i}}$ is an orthonormal basis of $M_{h}(\lambda_{i})$, for $\varphi\in M_{h}(\lambda_{i})$ with $\|\varphi\|_{a}=1$, there exists $\{\alpha_{j}\}$ satisfying $\varphi=\sum_{j=1}^{d_{i}}\alpha_{j}u^{h,o,0}_{ij}$ and $\sum_{j=1}^{d_{i}}\alpha_{j}^{2}=1$. Note that
$\operatorname{dist}_{a}\left(\varphi,U_{0}^{(i)}\right)=\left\|\left(\operatorname{I}-\mathcal{P}_{U_{0}^{(i)}}\right)\varphi\right\|_{a}\leqslant\sum_{j=1}^{d_{i}}|\alpha_{j}|\left\|\left(\operatorname{I}-\mathcal{P}_{U_{0}^{(i)}}\right)u^{h,o,0}_{ij}\right\|_{a}$
$=$ $\sum_{j=1}^{d_{i}}|\alpha_{j}|\operatorname{dist}_{a}\left(u^{h,o,0}_{ij},U_{0}^{(i)}\right)\leqslant\sum_{j=1}^{d_{i}}|\alpha_{j}|\operatorname{dist}_{a}\left(u^{h,o,0}_{ij},u_{ij}^{(0)}\right)\leqslant\sqrt{d_{i}}\varepsilon_{0}.$
@@
(31) $\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{0}^{(i)}\right)\leqslant\sqrt{d_{i}}\varepsilon_{0}\leqslant\sqrt{D}\varepsilon_{0},\quad i=1,\ldots,q.$
@@
$\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{1/2}^{(i)}\right)\leqslant\frac{\sqrt{D}\left(\gamma+\zeta_{0}\right)\varepsilon_{0}}{\sqrt{(g-\gamma-\zeta_{0})^{2}(1-D\varepsilon_{0}^{2})+D\left(\gamma+\zeta_{0}\right)^{2}\varepsilon^{2}_{0}}},\quad i=1,\ldots,q.$
@@
$\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{1}\right)\leqslant\frac{\sqrt{DN}\left(\gamma+\zeta_{0}\right)\varepsilon_{0}}{\sqrt{(g-\gamma-\zeta_{0})^{2}(1-D\varepsilon_{0}^{2})+D\left(\gamma+\zeta_{0}\right)^{2}\varepsilon^{2}_{0}}}:\triangleq\xi_{1}.$
We apply Proposition 2.3 and Theorem 3.7 to Eq. 27 and obtain that there exists an orthonormal basis $\{u^{h,o,1}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})$ with $b(u^{h,o,1}_{ij},u^{h,o,1}_{kl})=\delta_{ik}\delta_{jl}$ such that
(32) $\operatorname{dist}_{a}\left(u^{h,o,1}_{ij},u_{ij}^{(1)}\right)\leqslant C\xi_{1},\quad\lambda_{ij}^{(1)}-\lambda^{h}_{ij}\leqslant\lambda^{h}_{q+1,1}\xi_{1}^{2},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $C$ is a constant that is independent of $U_{1}$, i.e. independent of the iteration.
Set $\varepsilon_{1}=C\xi_{1}$ and $\zeta_{1}=\lambda^{h}_{q+1,1}\xi_{1}^{2}$, we then have
@@
$\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n+1}\right)\leqslant$ $\xi_{n+1}=\frac{\sqrt{DN}\left(\gamma+\zeta_{n}\right)\varepsilon_{n}}{\sqrt{(g-\gamma-\zeta_{n})^{2}(1-D\varepsilon_{n}^{2})+D\left(\gamma+\zeta_{n}\right)^{2}\varepsilon^{2}_{n}}},$
@@
$\operatorname{dist}_{a}\left(u^{h,o,n+1}_{ij},u_{ij}^{(n+1)}\right)\leqslant\varepsilon_{n+1}=$ $C\xi_{n+1}=\frac{C\sqrt{DN}\left(\gamma+\zeta_{n}\right)\varepsilon_{n}}{\sqrt{(g-\gamma-\zeta_{n})^{2}(1-D\varepsilon_{n}^{2})+D\left(\gamma+\zeta_{n}\right)^{2}\varepsilon^{2}_{n}}},$
for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$, where the constant $C$ is the same as the one in Eq. 32 due to the independence of iterations.

### v2

@@ Appendix A Detailed Proofs
For $\psi\in\bigoplus_{i=1}^{q}M_{h}(\lambda_{i})$ with $\|\psi\|=1$ and the orthonormal basis $\{v^{h}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})$ with $a(v^{h}_{ij},v^{h}_{kl})=\delta_{ik}\delta_{jl}$, there exists $\{\alpha_{ij}\}$ satisfying $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\alpha_{ij}^{2}=1$ and $\psi=\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\alpha_{ij}v^{h}_{ij}.$ It holds that
(25) $\operatorname{dist}\left(\psi,U_{n+1}\right)=\left\|\left(\operatorname{I}-\mathcal{P}_{U_{n+1}}\right)\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\alpha_{ij}v^{h}_{ij}\right\|$
$\leqslant$ $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}|\alpha_{ij}|\left\|\left(\operatorname{I}-\mathcal{P}_{U_{n+1}}\right)v^{h}_{ij}\right\|\leqslant\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\left\|\left(\operatorname{I}-\mathcal{P}_{U_{n+1}}\right)v^{h}_{ij}\right\|^{2}}$
$=$ $\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\operatorname{dist}^{2}\left(v^{h}_{ij},U_{n+1}\right)}\leqslant\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\operatorname{dist}^{2}\left(M_{h}(\lambda_{i}),U_{n+1}\right)}$
$\leqslant$ $\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\operatorname{dist}^{2}\left(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}\right)}\leqslant\sqrt{N}\max_{1\leqslant i\leqslant q}\operatorname{dist}(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}),$
@@
(26) $\operatorname{dist}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n+1}\right)\leqslant\sqrt{N}\max_{1\leqslant i\leqslant q}\operatorname{dist}(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)})\ll 1.$
It is clear that $U_{n+1}$ is a finite dimensional subspace of $V^{h}$. We apply Proposition 2.2 to Eq. 8 and obtain that $\lambda^{(n+1)}_{ij}\leqslant\lambda_{i+1,1}^{h}$ due to $\operatorname{dist}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n+1}\right)\ll 1$ for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$. Hence, it follows from Eq. 26, Proposition 2.2 and Theorem 3.7 applied to Eq. 8 that
$|\lambda_{ij}^{(n+1)}-\lambda^{h}_{ij}|\leqslant$ $\lambda_{i+1,1}^{h}\operatorname{dist}^{2}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n+1}\right)\leqslant\lambda_{i+1,1}^{h}\sqrt{N}\max_{1\leqslant i\leqslant q}\operatorname{dist}^{2}(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}),$
$\operatorname{dist}(u^{h,o}_{ij},u_{ij}^{(n+1)})\leqslant$ $\tilde{C}_{**}\operatorname{dist}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n+1}\right)\leqslant\tilde{C}_{**}\sqrt{N}\max_{1\leqslant i\leqslant q}\operatorname{dist}(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}),$
where the constant $\tilde{C}_{**}$ is independent of $U_{n+1/2}$ and may depend on $V^{h}$ (see Proposition 2.3 and Remark 2.4).
A.4 Proof of Proposition 4.2
Proof A.4.
@@
A.5 Proof of Theorem 4.3
Proof A.5.
@@
A simple calculation and the equation
$a(\mathcal{P}_{M_{h}(\lambda_{i})}v_{i}^{(0)},v)=a(v_{i}^{(0)},v),\quad v\in M_{h}(\lambda_{i})$
@@
We obtain from Eq. 2 and Eq. 11 that there exists $\varepsilon_{0}\in(0,1)$ such that
$\varepsilon_{0}\geqslant$ $\operatorname{dist}\left(U_{0}^{(i)},M_{h}(\lambda_{i})\right)\geqslant\operatorname{dist}\left(v_{i}^{(0)},M_{h}(\lambda_{i})\right)$
$=$ $\operatorname{dist}\left(v_{i}^{(0)},\mathcal{P}_{M_{h}(\lambda_{i})}v_{i}^{(0)}\right)=\frac{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\alpha_{rt}^{(i)}u^{h}_{rt}-\sum_{t=1}^{d_{i}}\alpha_{it}^{(i)}u^{h}_{it}\right\|}{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\alpha_{rt}^{(i)}u^{h}_{rt}\right\|},$
which yields
@@
(29) $\operatorname{dist}\left(v_{i}^{(1/2)},M_{h}(\lambda_{i})\right)=$ $\frac{\left\|v_{i}^{(1/2)}-\mathcal{P}_{M_{h}(\lambda_{i})}v_{i}^{(1/2)}\right\|}{\left\|v_{i}^{(1/2)}\right\|}$
@@
$\leqslant$ $\frac{\delta_{0}\varepsilon_{0}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{0}^{2})+\delta_{0}^{2}\varepsilon^{2}_{0}}}:=\varepsilon_{1},\quad i=1,\ldots,q.$
@@
$\operatorname{dist}\left(M_{h}(\lambda_{i}),U_{1/2}^{(i)}\right)\leqslant\varepsilon_{1},\quad i=1,2,\ldots,q.$
@@
$\operatorname{dist}\left(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}\right)\leqslant\varepsilon_{n+1}=\frac{\delta_{0}\varepsilon_{n}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{n}^{2})+\delta_{0}^{2}\varepsilon^{2}_{n}}},\quad\forall n\in\mathbb{N}.$
@@
A.6 Proof of Theorem 4.4
Proof A.6.
@@
Since $\{u^{h,o,0}_{ij}\}_{i=1}^{d_{i}}$ is an orthonormal basis of $M_{h}(\lambda_{i})$, for $\varphi\in M_{h}(\lambda_{i})$ with $\|\varphi\|=1$, there exists $\{\alpha_{j}\}$ satisfying $\varphi=\sum_{j=1}^{d_{i}}\alpha_{j}u^{h,o,0}_{ij}$ and $\sum_{j=1}^{d_{i}}\alpha_{j}^{2}=1$. Note that
$\operatorname{dist}\left(\varphi,U_{0}^{(i)}\right)=\left\|\left(\operatorname{I}-\mathcal{P}_{U_{0}^{(i)}}\right)\varphi\right\|\leqslant\sum_{j=1}^{d_{i}}|\alpha_{j}|\left\|\left(\operatorname{I}-\mathcal{P}_{U_{0}^{(i)}}\right)u^{h,o,0}_{ij}\right\|$
$=$ $\sum_{j=1}^{d_{i}}|\alpha_{j}|\operatorname{dist}\left(u^{h,o,0}_{ij},U_{0}^{(i)}\right)\leqslant\sum_{j=1}^{d_{i}}|\alpha_{j}|\operatorname{dist}\left(u^{h,o,0}_{ij},u_{ij}^{(0)}\right)\leqslant\sqrt{d_{i}}\varepsilon_{0}.$
@@
(31) $\operatorname{dist}\left(M_{h}(\lambda_{i}),U_{0}^{(i)}\right)\leqslant\sqrt{d_{i}}\varepsilon_{0}\leqslant\sqrt{D}\varepsilon_{0},\quad i=1,\ldots,q.$
@@
$\operatorname{dist}\left(M_{h}(\lambda_{i}),U_{1/2}^{(i)}\right)\leqslant\frac{\sqrt{D}\left(\gamma+\zeta_{0}\right)\varepsilon_{0}}{\sqrt{(g-\gamma-\zeta_{0})^{2}(1-D\varepsilon_{0}^{2})+D\left(\gamma+\zeta_{0}\right)^{2}\varepsilon^{2}_{0}}},\quad i=1,\ldots,q.$
@@
$\operatorname{dist}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{1}\right)\leqslant\frac{\sqrt{DN}\left(\gamma+\zeta_{0}\right)\varepsilon_{0}}{\sqrt{(g-\gamma-\zeta_{0})^{2}(1-D\varepsilon_{0}^{2})+D\left(\gamma+\zeta_{0}\right)^{2}\varepsilon^{2}_{0}}}:=\xi_{1}.$
We apply Theorem 4.1 to Eq. 13 and obtain that there exists an orthonormal basis $\{u^{h,o,1}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})$ with $b(u^{h,o,1}_{ij},u^{h,o,1}_{kl})=\delta_{ik}\delta_{jl}$ such that
(32) $\operatorname{dist}\left(u^{h,o,1}_{ij},u_{ij}^{(1)}\right)\leqslant\tilde{C}_{**}\xi_{1},\quad\lambda_{ij}^{(1)}-\lambda^{h}_{ij}\leqslant\lambda^{h}_{q+1,1}\xi_{1}^{2},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $\tilde{C}_{**}$ is a constant that is independent of $U_{1}$, i.e., independent of the iteration.
Set $\varepsilon_{1}=\tilde{C}_{**}\xi_{1}$ and $\zeta_{1}=\lambda^{h}_{q+1,1}\xi_{1}^{2}$, we then have
@@
$\operatorname{dist}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n+1}\right)\leqslant$ $\xi_{n+1}=\frac{\sqrt{DN}\left(\gamma+\zeta_{n}\right)\varepsilon_{n}}{\sqrt{(g-\gamma-\zeta_{n})^{2}(1-D\varepsilon_{n}^{2})+D\left(\gamma+\zeta_{n}\right)^{2}\varepsilon^{2}_{n}}},$
@@
$\operatorname{dist}\left(u^{h,o,n+1}_{ij},u_{ij}^{(n+1)}\right)\leqslant\varepsilon_{n+1}=$ $\tilde{C}_{**}\xi_{n+1}=\frac{\tilde{C}_{**}\sqrt{DN}\left(\gamma+\zeta_{n}\right)\varepsilon_{n}}{\sqrt{(g-\gamma-\zeta_{n})^{2}(1-D\varepsilon_{n}^{2})+D\left(\gamma+\zeta_{n}\right)^{2}\varepsilon^{2}_{n}}},$
for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$, where the constant $\tilde{C}_{**}$ is the same as the one in Eq. 32 due to the independence of iterations.